OpenStudy (anonymous):

what is the next number in the geometric sequence -3.42,10.26,-30.78,92.34

jimthompson5910 (jim_thompson5910):

Notice how \[\Large \frac{10.26}{-3.42} = -3\] \[\Large \frac{-30.78}{10.26} = -3\] \[\Large \frac{92.34}{-30.78} = -3\] So you're multiplying each term by -3 to get the next term.